suppose a consumer consuming two comodities X and Y has the following utility function U=X^0.4Y^0.6. if the price of good X and Y are 2 and 3 respectively and the income constraint is birr 50. find the quantities of X anf Y which maximize utility


1
Expert's answer
2021-10-10T16:31:10-0400

"U(X,Y) = X^{0.4}Y^{0.6}"

Budget line

"50 = X+Y \\

P(X)=2 \\

P(Y) = 3"

At optimum : "\frac{MU(X)}{P(X)}= \frac{MU(Y)}{P(Y)}"

"MU(Y) = \frac{ΔU}{ΔX} = 0.4X^{-0.6}Y^{0.6} \\

MU(y) = \frac{ΔU}{ΔY} = 0.6X^{0.4}Y^{-0.4} \\

\frac{0.4X^{-0.6}Y^{0.6}}{2} = \frac{0.6X^{0.4}Y^{-0.4}}{3} \\

0.2X^{-0.6}Y^{0.6}=0.2X^{0.4}Y^{-0.4} \\

X=Y"

Putting X=Y in budget line

"50 = Y+Y \\

50 = 2Y \\

Y = \frac{50}{2} = 25 \\

X= 25 \\

U(X,Y)_{max} = 25^{0.4} \times 25^{0.6} = 25"